Please help me solve this permutation: "A men's department store sells 3 different suit jackets, 6 different shirts, 8 different ties, and 4 different pairs of pants. How many different suits consisting of a jacket, shirt, tie, and pants are possible?"
Found 2 solutions by hokies, palanisamy:
Answer by hokies(65)

Multiply them all together:
Combinations = 3*6*8*4 = 576

Answer by palanisamy(496)

Given,
A men's department store sells 3 different suit jackets, 6 different shirts, 8 different ties, and 4 different pairs of pants.
Out of 3 different suit jackets one jacket can be selected in 3 ways.
Out of 6 different shirts one shirt can be selected in 6 ways.
Out of 8 different ties one tie can be selected in 8 ways.
Out of 4 different pairs of pants one can be selected in 4 ways.
Total number of ways of selection = 3*6*8*4 = 576